# Regression-Based Single-Point Zeroth-Order Optimization

### v1

@@ Abstract
Zeroth-order optimization (ZO) is widely used for solving black-box optimization and control problems. In particular, single-point ZO (SZO) is well-suited to online or dynamic problem settings due to its requirement of only a single function evaluation per iteration. However, SZO suffers from high gradient estimation variance and slow convergence, which severely limit its practical applicability. Despite recent advances, the convergence of existing SZO methods remains inferior to that of two-point ZO methods. To overcome this limitation, we propose a novel yet simple SZO framework, termed regression-based SZO (RESZO), which substantially enhances the convergence rate. Unlike conventional ZO methods that rely solely on function evaluations at the current point for gradient estimation, RESZO improves gradient estimation by effectively leveraging historical function evaluations from previous iterations. Specifically, RESZO constructs a surrogate function via regression using recent historical evaluations and employs the gradient of this surrogate function for iterative updates. Two variants of RESZO, which fit linear and quadratic surrogate functions respectively, are introduced. Theoretically, we provide a non-asymptotic convergence analysis for the linear variant of RESZO, showing that its convergence rates are comparable to those of two-point ZO methods for both smooth nonconvex and strongly convex functions. Moreover, extensive numerical experiments demonstrate that RESZO not only matches but empirically outperforms two-point ZO in terms of function query complexity.
@@ 1 Introduction
where $\bm{x}\in\mathbb{R}^{d}$ is the decision variable of dimension $d$. The objective function $f:\mathbb{R}^{d}\to\mathbb{R}$ is unknown, and only its zeroth-order information, namely function evaluations, is accessible. As a result, the first-order gradients of function $f$ are unavailable, and problem (1) can be solved solely based on function evaluations. This type of black-box optimization problems has attracted considerable recent attention and covers a broad range of applications, including power system control [1, 2], reinforcement learning [3, 4], neural network training [5, 6], sensor management [7], etc.
@@
Motivation. At each iteration $t$, classic ZO methods estimate the gradient $\nabla f(\bm{x}_{t})$ solely based on the function evaluations at the current point $\bm{x}_{t}$, while all historical function evaluations obtained from previous iterations are discarded. However, these past function evaluations contain valuable information about the unknown function and can help with the gradient estimations for subsequent iterations, especially those obtained in the most recent iterations. Moreover, in many practical applications, such as clinical trials, real-world experiments, and long-term complex simulations, obtaining a single function evaluation can be costly and laborious, making it crucial to fully exploit all available evaluations to enhance efficiency. Then, a natural question is
@@
Contribution. Motivated by this question, we propose a novel yet simple single-point ZO framework, termed Regression-based Single-point Zeroth-order Optimization (RESZO). In contrast to classic ZO methods that rely solely on function evaluations at the current point for gradient estimation, the proposed RESZO method significantly enhances gradient estimation by leveraging available historical function evaluations from previous iterations. Specifically, the RESZO method queries the function value only once at each iteration, and fits a surrogate model based on recent historical function evaluations via regression. Then, the surrogate model’s gradient is used as the estimated gradient for iterative updates. We introduce two variants of the RESZO method, termed L-RESZO and Q-RESZO, which fit linear and quadratic surrogate functions, respectively. A key ingredient of the proposed RESZO method is adding random perturbations to function evaluations to ensure sufficient exploration. Theoretically (see Table 1), we show that for both smooth nonconvex and strongly convex functions, the linear variant of our algorithm (L-RESZO) achieves convergence rates that are only worse than those of two-point ZO methods by a factor $C_{d}>0$, where $C_{d}$ is a parameter defined in (21) in Assumption 2. This represents a substantial improvement over the rates achieved by existing SZO methods (see Table 1). While theoretically bounding the term $C_{d}$ is challenging, strong empirical evidence shows that $C_{d}$ scales as $\mathcal{O}(\sqrt{d})$ under appropriate algorithmic settings; see Section 4 for more details. To the best of our knowledge, this is the first work to present single-point ZO methods whose convergence rates match those of two-point ZO methods. Moreover, extensive numerical experiments demonstrate that RESZO not only matches but empirically outperforms two-point ZO in terms of function query complexity.
@@
Related Work. The classic SZO method, formulated in (2), is first proposed by [9], which estimates gradients by querying only a single function value per iteration. In [13], a self-concordant barrier function is used to enhance the convergence of SZO for solving bandit smooth convex optimization. By leveraging high-order smoothness, reference [16] derives an improved complexity bound for SZO in the context of stochastic convex optimization. To further enhance the convergence of SZO, reference [15] proposes a residual-feedback SZO method, formulated in (2), which reuses the function evaluation from the last iteration for gradient estimation, resulting in significant performance improvement. Additionally, reference [11] leverages the idea of high-pass and low-pass filters from extremum seeking control to accelerate SZO convergence. Notably, the integration of a high-pass filter coincides with the residual-feedback approach in [15]. This paper follows a similar direction to [11, 15]; however, rather than reusing only the most recent function evaluation, we leverage many historical function evaluations from previous iterations to fit surrogate models. In addition, our RESZO method is related to the linear regression and interpolation approaches [17, 18, 19] used in derive-free optimization (DFO) [20]. The fundamental difference is that these methods [17, 18, 19] perform multiple function evaluations near the current point for linear interpolation or regression. In contrast, our approach queries the function value only once at each iteration and fits surrogate models by fully recycling historical function evaluations from past iterations.
2 Preliminaries on Zeroth-Order Optimization
@@
[…]
where $\delta\!>\!0$ is a scalar parameter called smoothing radius, $\eta\!>\!0$ is the step size, and $\bm{u}_{t}\in\mathbb{R}^{d}$ is an independent identically distributed random direction vector sampled from $\mathrm{Unif}(\mathbb{S}_{d-1})$, namely the uniform distribution on the unit sphere $\mathbb{S}_{d-1}$, for all $k=0,1,2,\cdots$. The random direction vector $\bm{u}_{t}$ can also be sampled from the multivariate Gaussian distribution $\mathcal{N}(\mathbf{0},\bm{I})$ or other zero-mean distributions. In (2), $\mathsf{G}^{1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ denotes the SZO gradient estimator of function $f$.
Recent works [15, 11] have introduced a new variant of SZO, referred to as residual-feedback SZO (RSZO) [15] or high-pass filter-integrated SZO [11], which is given by (2):
[…]
$\qquad\bm{x}_{t+1}=\bm{x}_{t}-\eta\cdot\underbrace{\frac{d}{\delta}\Big{(}f(\bm{x}_{t}+\delta\bm{u}_{t})-f(\bm{x}_{t-1}+\delta\bm{u}_{t-1})\Big{)}\bm{u}_{t}}_{\coloneq\,\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)},\qquad t=1,2,\cdots.$ (3)
At each iteration $t$, the gradient estimator $\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ of the RSZO method (2) reuses the function evaluation $f(\bm{x}_{t\!-\!1}\!+\!r\bm{u}_{t\!-\!1})$ queried at the last iteration $t\!-\!1$. This approach substantially reduces gradient estimation variance and improves convergence compared with SZO (2).
@@
[…]
$\qquad\qquad\bm{x}_{t+1}=\bm{x}_{t}-\eta\cdot\underbrace{\frac{d}{2\delta}\Big{(}f(\bm{x}_{t}+\delta\bm{u}_{t})-f(\bm{x}_{t}-\delta\bm{u}_{t})\Big{)}\bm{u}_{t}}_{\coloneq\,\mathsf{G}^{2}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)},\quad t=0,1,2,\cdots,$ (4)
@@
$\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big{[}\mathsf{G}^{1}_{f}(\bm{x};\bm{u},\delta)\big{]}=\,$ $\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big{[}\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x};\bm{u},\delta)\big{]}$
=\,= $\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big{[}\mathsf{G}^{2}_{f}(\bm{x};\bm{u},\delta)\big{]}=\nabla\hat{f}(\bm{x};\delta),$ (5)
where $\hat{f}(\bm{x};\delta):=\mathbb{E}_{\bm{v}\sim\mathrm{Unif}(\mathbb{B}_{d})}\big{[}f(\bm{x}+\delta\bm{v})\big{]}$ is a smoothed version of the original function $f$ [9]. Despite having the same expectation, the two-point ZO gradient estimator $\mathsf{G}^{2}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ exhibits lower variance and achieves faster convergence, compared with the single-point ZO gradient estimators $\mathsf{G}^{1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ and $\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$. See [8, 11] for more details.
@@ 3 Regression-Based Single-Point Zeroth-Order Optimization
In this paper, we focus on the class of SZO methods that require only one function evaluation per iteration. Below, we introduce the proposed RESZO method and two specific variants, L-RESZO and Q-RESZO, which fit linear and quadratic surrogate models using historical function evaluations.
@@ 3.1 Linear Regression-Based Single-Point Zeroth-Order Optimization (L-RESZO)
At each iteration $t$, the available function evaluations from all previous iterations are given by $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ with $\hat{\bm{x}}_{k}=\bm{x}_{k}+\delta\bm{u}_{k}$. Here, $\bm{u}_{k}\in\mathbb{R}^{d}$ is a random direction vector sampled from the uniform distribution $\mathrm{Unif}(\mathbb{S}_{d-1})$ over the unit sphere $\mathbb{S}_{d-1}$, or from the multivariate Gaussian distribution $\mathcal{N}(\bm{0},\bm{I})$; and $\delta>0$ is a scalar parameter called smoothing radius. These are consistent with the classic ZO methods (2)-(4). Note that function evaluations are performed at the perturbed points $\hat{\bm{x}}_{k}$ rather than the iterate points ${\bm{x}}_{k}$ to promote sufficient exploration of the optimization space. The necessity of introducing such perturbations is further illustrated by Remark 1 and the numerical experiments in Section 5.2.
@@
Based on the first-order Taylor expansion at the current perturbed point $\hat{\bm{x}}_{t}$, the available function evaluations $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ can be expressed as (6):
$\qquad\qquad f(\hat{\bm{x}}_{k})=f(\hat{\bm{x}}_{t})+\nabla f(\hat{\bm{x}}_{t})^{\top}(\hat{\bm{x}}_{k}-\hat{\bm{x}}_{t})+\epsilon_{t,k},\qquad k=0,1,\cdots,t,$ (6)
@@
where $\epsilon_{t,k}$ denotes the corresponding high-order residual term. Accordingly, we propose to fit a linear surrogate function $f^{\mathrm{s1}}_{t}(\bm{x})$ in (7) to locally approximate the original function $f(\bm{x})$ in (1):
$f^{\mathrm{s1}}_{t}(\bm{x})\coloneq f(\hat{\bm{x}}_{t})+\bm{g}_{t}^{\top}({\bm{x}}-\hat{\bm{x}}_{t})+c_{t},$ (7)
@@
where the linear coefficient vector $\bm{g}_{t}\in\mathbb{R}^{d}$ and the intercept $c_{t}\in\mathbb{R}$ can be obtained through linear regression on the available sample points $\{\hat{\bm{x}}_{k},f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$. Then, the gradient of the surrogate function $\nabla f^{\mathrm{s1}}_{t}(\bm{x})=\bm{g}_{t}$ can be used as the estimator for the true gradient $\nabla f(\bm{x})$.
@@
To ensure that the fitted surrogate model $f^{\mathrm{s1}}_{t}(\bm{x})$ remains up-to-date with the most recent function evaluations, we employ the recursive linear regression in a sliding window, which uses only the latest $m$ function evaluations $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$ to fit the linear model at each iteration $t$. Specifically, at each iteration $t$, we solve the following least squares problem (8) to obtain $\bm{g}_{t}$ and $c_{t}$:
$\min_{\bm{g}_{t},c_{t}}\,\sum_{i=1}^{m}\Big{[}f^{\mathrm{s1}}_{t}(\hat{\bm{x}}_{t-i+1})-f(\hat{\bm{x}}_{t-i+1})\Big{]}^{2}\ \Longrightarrow\ \begin{bmatrix}\bm{g}_{t}\\ c_{t}\end{bmatrix}=(X_{t}^{\top}X_{t})^{\dagger}X_{t}^{\top}\bm{y}_{t},$ (8)
@@
where $\dagger$ denotes the Moore–Penrose pseudoinverse, $\bm{\Delta}^{x}_{t,i}\!\coloneq\!\hat{\bm{x}}_{t-i+1}\!-\!\hat{\bm{x}}_{t},\Delta^{f}_{t,i}\!\coloneq\!f(\hat{\bm{x}}_{t-i+1})\!-\!f(\hat{\bm{x}}_{t})$,
$X_{t}\coloneq\begin{bmatrix}{\bm{\Delta}^{x}_{t,1}}^{\top}&1\\ {\bm{\Delta}^{x}_{t,2}}^{\top}&1\\ \vdots&\vdots\\ {\bm{\Delta}^{x}_{t,m}}^{\top}&1\end{bmatrix}\in\mathbb{R}^{m\times(d+1)},\quad\bm{y}_{t}\coloneq\begin{bmatrix}\Delta^{f}_{t,1}\\ \Delta^{f}_{t,2}\\ \vdots\\ \Delta^{f}_{t,m}\end{bmatrix}\in\mathbb{R}^{m}.$ (9)
@@
Consequently, the proposed Linear Regression-based Single-point Zeroth-order Optimization (L-RESZO) method for solving problem (1) is presented as Algorithm 1. Since the linear regression requires $m$ function evaluations, we initially employ a classical ZO algorithm, such as the RSZO method (2), to perform iterative updates and accumulate sufficient sample points, as shown in Step 3. After $m$ iterations, the algorithm switches to using the linear regression-based gradient estimator $\bm{g}_{t}$ for subsequent iterative updates, as shown in Steps 6-8.
@@
1:  Inputs: Initial point $\bm{x}_{0}$, stepsizes $\tilde{\eta},\eta$, smoothing radius values $\tilde{\delta},\delta$, sample size $m$, and total iteration number $T$.
2:  for $t=0:m-1$ do
3:     Perform the RSZO method (2) with $\tilde{\eta},\tilde{\delta}$ for iterative updates.
@@
5:  for $t=m:T$ do
6:     Randomly sample $\bm{u}_{t}\sim\mathrm{Unif}(\mathbb{S}_{d-1})$; perform a function evaluation on the perturbed value $\hat{\bm{x}}_{t}=\bm{x}_{t}+\delta\bm{u}_{t}$ to obtain $f(\hat{\bm{x}}_{t})$; collect $m$ latest sample points $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$.
7:     Solve for the linear coefficient $\bm{g}_{t}$ via (8).
8:     Update with $\bm{x}_{t+1}=\bm{x}_{t}-\eta\cdot\bm{g}_{t}$.
9:  end for
@@
In addition, the sample size $m$ is typically chosen to be comparable to the dimension $d$ of $\bm{x}$, or slightly larger, to ensure that $X_{t}$ is of full column rank and the matrix $X_{t}^{\top}X_{t}$ in (8) is invertible. In this case, one can replace the pseudoinverse $(\cdot)^{\dagger}$ in (8) by the regular matrix inverse $(\cdot)^{-1}$.

### v2

# Accelerating Single-Point Zeroth-Order Optimization with Regression-Based Gradient Surrogates

@@ Abstract
Zeroth-order optimization (ZO) is widely used for solving black-box optimization and control problems. In particular, single-point ZO (SZO) is well-suited to online or dynamic problem settings due to its requirement of only a single function evaluation per iteration. However, SZO suffers from high gradient estimation variance and slow convergence, which severely limit its practical applicability. To overcome this limitation, we propose a novel yet simple SZO framework termed regression-based SZO (ReSZO), which substantially enhances the convergence rate. Specifically, ReSZO constructs a surrogate function via regression using historical function evaluations and employs the gradient of this surrogate function for iterative updates. Two instantiations of ReSZO, which fit linear and quadratic surrogate functions respectively, are introduced. Moreover, we provide a non-asymptotic convergence analysis for the linear instantiation of ReSZO, showing that its convergence rates are comparable to those of two-point ZO methods. Extensive numerical experiments demonstrate that ReSZO empirically converges two to three times faster than two-point ZO in terms of function query complexity.
@@ 1 Introduction
where $\bm{x}\in\mathbb{R}^{d}$ is the decision variable of dimension $d$. The objective function $f:\mathbb{R}^{d}\to\mathbb{R}$ is unknown, and only its zeroth-order information, namely function evaluations, is accessible. As a result, the first-order gradients of function $f$ are unavailable, and problem (1) can be solved solely based on function evaluations. This black-box optimization has attracted considerable recent attention and covers a broad range of applications, including power system control [4, 8], reinforcement learning [18, 16], neural network training [2, 3], sensor management [17], etc.
@@
Motivation. At each iteration $t$, classic ZO methods estimate the gradient $\nabla f(\bm{x}_{t})$ solely based on the function evaluations at the current iterate $\bm{x}_{t}$, while all historical function evaluations obtained from previous iterations are discarded. However, these past function evaluations contain valuable information about the unknown function and can help with the gradient estimations for subsequent iterations, especially those obtained in the most recent iterations. Moreover, in many practical applications, e.g., clinical trials, real-world experiments, and long-horizon simulations, obtaining even a single function evaluation can be costly and laborious, making it crucial to fully exploit all available evaluations to improve efficiency. Then, a natural question is
@@
Contribution. Motivated by this question, we propose a novel yet simple single-point ZO framework termed Regression-based Single-point Zeroth-order Optimization (ReSZO). Unlike classical ZO methods that estimate gradients using only function evaluations at the current iterate, ReSZO exploits recent historical function evaluations to significantly improve gradient estimation while requiring only one function query per iteration. Our main contributions are summarized as follows:
  • 1)
    (New Framework and Algorithms). We introduce the ReSZO framework and develop two concrete instantiations, namely L-ReSZO and Q-ReSZO, which construct linear and quadratic surrogate models, respectively. By leveraging historical function evaluations together with an adaptive perturbation scheme, the proposed methods substantially improve convergence efficiency over existing SZO methods and demonstrate strong empirical performance.
  • 2)
    (Theoretical Analysis). We provide theoretical analysis for the convergence of ReSZO. Specifically, for both smooth nonconvex and strongly convex objectives, we show that under certain conditions, L-ReSZO achieves convergence rates that are within a factor $C_{d}\!>\!0$ of those attained by two-point ZO methods (see Table 1). While tightly bounding $C_{d}$ is challenging, we offer theoretical justification that $C_{d}$ scales as $\mathcal{O}(d)$. We also provide extensive empirical evidence indicating that, under suitable algorithmic settings, the practical scaling follows $C_{d}=\mathcal{O}(\sqrt{d})$ (see Section 4).
  • 3)
    (Empirical Evaluation). We conduct extensive numerical experiments on a variety of benchmark problems, demonstrating that ReSZO not only matches but empirically converges two to three times faster than two-point ZO in terms of function query complexity.
@@
Related Work. The classic SZO method, formulated in (2), is first proposed by [12], which estimates gradients by querying only a single function value per iteration. In [10], a self-concordant barrier function is used to enhance the convergence of SZO for solving bandit smooth convex optimization. By leveraging high-order smoothness, reference [20] derives an improved complexity bound for SZO in the context of stochastic convex optimization. To further enhance the convergence of SZO, reference [27] proposes a residual-feedback SZO method, formulated in (2), which reuses the function evaluation from the last iteration for gradient estimation, resulting in significant performance improvement. Additionally, reference [7] leverages the idea of high-pass and low-pass filters from extremum seeking control to accelerate SZO convergence. Notably, the integration of a high-pass filter coincides with the residual-feedback approach in [27]. This paper follows a similar direction to [7, 27]; however, rather than reusing only the most recent function evaluation, we leverage many historical function evaluations from previous iterations to fit surrogate models. In addition, our ReSZO method is related to the linear regression and interpolation approaches [9, 1, 25] used in derive-free optimization (DFO) [15]. The fundamental difference is that these methods [9, 1, 25] perform multiple function evaluations near the current iterate for linear interpolation or regression. In contrast, our method queries the function value only once at each iteration and fits surrogate models by recycling historical function evaluations. We also note a recent SZO algorithm [14] that achieves strong convergence rates. However, unlike our approach, it assumes access to an oracle capable of handling complex-valued queries, which limits its applicability in standard zeroth-order settings where the oracle takes only real-valued queries.
2 Preliminaries
@@
$\ \textbf{Single-point ZO (SZO)}:\ \bm{x}_{t+1}=\bm{x}_{t}-\eta\cdot\underbrace{\frac{d}{\delta}f(\bm{x}_{t}+\delta\bm{u}_{t})\bm{u}_{t}}_{:=\,\mathsf{G}^{1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)},\quad t=0,1,2,\cdots$ (2)
where $\delta\!>\!0$ is a scalar parameter called smoothing radius, $\eta\!>\!0$ is the step size, and $\bm{u}_{t}\in\mathbb{R}^{d}$ is an independent identically distributed random direction vector sampled from $\mathrm{Unif}(\mathbb{S}_{d-1})$, namely the uniform distribution on the unit sphere $\mathbb{S}_{d-1}$, for all $t=0,1,2,\cdots$. The random direction vector $\bm{u}_{t}$ can also be sampled from the multivariate Gaussian distribution $\mathcal{N}(\mathbf{0},\bm{I})$ or other zero-mean distributions. In (2), $\mathsf{G}^{1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ denotes the SZO gradient estimator of function $f$.
Recent works have introduced a new variant of SZO, referred to as residual-feedback SZO (RSZO) [27] or high-pass filter-integrated SZO [7], which is given by (2):
$\ \textbf{Residual-feedback SZO (RSZO)}:\text{for }t=1,2,\cdots,$
$\ \bm{x}_{t+1}=\bm{x}_{t}\!-\!\eta\cdot\underbrace{\frac{d}{\delta}\Big(f(\bm{x}_{t}\!+\!\delta\bm{u}_{t})\!-\!f(\bm{x}_{t-1}\!+\!\delta\bm{u}_{t-1})\Big)\bm{u}_{t}}_{\coloneq\,\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)}.$ (3)
At each iteration $t$, the gradient estimator $\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ of the RSZO method (2) reuses the function evaluation $f(\bm{x}_{t\!-\!1}\!+\!r\bm{u}_{t\!-\!1})$ queried at the last iteration t1t\!-\!1. This approach substantially reduces gradient estimation variance and improves convergence compared with SZO (2).
@@
$\ \textbf{Two-point ZO (TZO)}:\text{ for }t=0,1,2,\cdots$
$\ \bm{x}_{t+1}=\bm{x}_{t}-\eta\cdot\underbrace{\frac{d}{2\delta}\Big(f(\bm{x}_{t}+\delta\bm{u}_{t})-f(\bm{x}_{t}-\delta\bm{u}_{t})\Big)\bm{u}_{t}}_{\coloneq\,\mathsf{G}^{2}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)},$ (4)
@@
$\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big[\mathsf{G}^{1}_{f}(\bm{x};\bm{u},\delta)\big]$
$=\,$ $\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big[\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x};\bm{u},\delta)\big]$
$=\,$ $\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big[\mathsf{G}^{2}_{f}(\bm{x};\bm{u},\delta)\big]=\nabla\hat{f}(\bm{x};\delta),$ (5)
where $\hat{f}(\bm{x};\delta):=\mathbb{E}_{\bm{v}\sim\mathrm{Unif}(\mathbb{B}_{d})}\big[f(\bm{x}+\delta\bm{v})\big]$ is a smoothed version of function $f$ [12]. Despite having the same expectation, the two-point ZO gradient estimator $\mathsf{G}^{2}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ exhibits lower variance and achieves faster convergence, compared with the single-point ZO gradient estimators $\mathsf{G}^{1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ and $\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$. See [19, 7] for more details.
@@ 3 Regression-Based Single-Point Zeroth-Order Optimization
In this paper, we focus on the class of SZO methods that require only one function evaluation per iteration. Below, we introduce the ReSZO method and its two instantiations, L-ReSZO and Q-ReSZO, which fit linear and quadratic surrogate models using historical function evaluations.
At each iteration $t$, the available function evaluations from all previous iterations are given by $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ with $\hat{\bm{x}}_{k}=\bm{x}_{k}+\delta\bm{u}_{k}$. Here, $\bm{u}_{k}\in\mathbb{R}^{d}$ is a random direction vector sampled from the uniform distribution $\mathrm{Unif}(\mathbb{S}_{d-1})$ over the unit sphere $\mathbb{S}_{d-1}$, and $\delta>0$ is a scalar parameter called smoothing radius. These are the same as the classic ZO methods (2)-(4). Note that function evaluations are performed at the perturbed points $\hat{\bm{x}}_{k}$ rather than the iterate points ${\bm{x}}_{k}$ to promote exploration of the optimization space. The necessity of introducing such perturbations is further illustrated by Remark 1 and the numerical experiments in Section 5.2.
@@ 3.1 Linear Regression-Based Single-Point Zeroth-Order Optimization (L-ReSZO)
Based on the first-order Taylor expansion at the current perturbed point $\hat{\bm{x}}_{t}$, the available function evaluations $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ can be expressed as (6): for $k=0,1\dots,t$,
$f(\hat{\bm{x}}_{k})\!=\!f(\hat{\bm{x}}_{t})\!+\!\nabla f(\hat{\bm{x}}_{t})^{\top}\big(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t}\big)\!+\!\epsilon_{t,k}(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t}),$ (6)
@@
where the function $\epsilon_{t,k}(\cdot)$ represents the corresponding high-order residual term. Accordingly, we propose to fit a linear surrogate function $f^{\mathrm{s1}}_{t}(\bm{x})$ in (7) to locally approximate the original function $f(\bm{x})$ in (1):
$f^{\mathrm{s1}}_{t}(\bm{x})\coloneq f(\hat{\bm{x}}_{t})+\bm{g}_{t}^{\top}\big({\bm{x}}-\hat{\bm{x}}_{t}\big),$ (7)
@@
where the linear coefficient vector $\bm{g}_{t}\in\mathbb{R}^{d}$ can be obtained through linear regression on the available sample points $\{\hat{\bm{x}}_{k},f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$. Then, the gradient of the surrogate function $\nabla f^{\mathrm{s1}}_{t}(\bm{x})=\bm{g}_{t}$ can be used as the estimator for the true gradient $\nabla f(\bm{x})$. A key feature of the linear surrogate function (7) is the omission of a constant intercept term, which enforces $f^{\mathrm{s1}}_{t}(\hat{\bm{x}}_{t})=f(\hat{\bm{x}}_{t})$ and reduces model bias.
@@
To ensure that the fitted surrogate model $f^{\mathrm{s1}}_{t}(\bm{x})$ remains up-to-date with the most recent function evaluations, we employ recursive linear regression in a sliding window, which uses only the latest $m$ function evaluations $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$ to fit the linear model at each iteration $t$. Specifically, at each iteration $t$, we solve the following least squares problem (8) to obtain $\bm{g}_{t}$:
$\ \min_{\bm{g}_{t}}\,\sum_{i=1}^{m}\Big[f^{\mathrm{s1}}_{t}(\hat{\bm{x}}_{t-i+1})-f(\hat{\bm{x}}_{t-i+1})\Big]^{2}\Longrightarrow\ \bm{g}_{t}=(X_{t}^{\top}X_{t})^{\dagger}X_{t}^{\top}\bm{y}_{t},$ (8)
@@
where $\dagger$ denotes the Moore–Penrose pseudoinverse, $\bm{\Delta}^{x}_{t,i}\!\coloneq\!\hat{\bm{x}}_{t-i}\!-\!\hat{\bm{x}}_{t},\Delta^{f}_{t,i}\!\coloneq\!f(\hat{\bm{x}}_{t-i})\!-\!f(\hat{\bm{x}}_{t})$,
$X_{t}\coloneq\!\!\begin{bmatrix}{\bm{\Delta}^{x}_{t,1}}^{\top}\\ {\bm{\Delta}^{x}_{t,2}}^{\top}\\ \vdots\\ {\bm{\Delta}^{x}_{t,m-1}}^{\top}\end{bmatrix}\!\!\in\mathbb{R}^{(m\!-\!1)\times d},\ \bm{y}_{t}\coloneq\!\!\begin{bmatrix}\Delta^{f}_{t,1}\\ \Delta^{f}_{t,2}\\ \vdots\\ \Delta^{f}_{t,m-1}\end{bmatrix}\!\!\in\mathbb{R}^{m\!-\!1}.$ (9)
@@
Consequently, the proposed Linear Regression-based Single-point Zeroth-order Optimization (L-ReSZO) method for solving problem (1) is presented as Algorithm 1. Since the method requires $m$ function evaluations for regression, we initially employ a classical ZO algorithm, e.g., the RSZO method (2), to perform iterative updates and accumulate sufficient sample points, as shown in Steps 2-4. After $m$ iterations, the algorithm switches to using the linear regression-based gradient estimator $\bm{g}_{t}$ for subsequent iterative updates, as shown in Steps 5-10. The sample size $m$ is typically chosen to be comparable to the dimension $d$ of $\bm{x}$, or slightly larger, to ensure that $X_{t}$ is of full column rank and the matrix $X_{t}^{\top}X_{t}$ in (8) is invertible. In this case, one can replace the pseudoinverse $(\cdot)^{\dagger}$ in (8) by the regular matrix inverse $(\cdot)^{-1}$.
@@
1:Input: Initial point $\bm{x}_{0}$; step sizes $\tilde{\eta},\eta$; smoothing radius $\tilde{\delta}$; sample size $m$; total iterations $T$.
2:for $t=0$ to m1m-1 do
3:  Perform RSZO update (i.e., Eq. (2)) with $\tilde{\eta},\tilde{\delta}$.
@@
5:for $t=m$ to T1T-1 do
6:  Use an adaptive smoothing radius by setting:
$\delta_{t}\!=\!\|\bm{x}_{t}\!-\!\bm{x}_{t-1}\|.$ (10)
7:  Sample $\bm{u}_{t}\sim\mathrm{Unif}(\mathbb{S}_{d-1})$; set $\hat{\bm{x}}_{t}=\bm{x}_{t}+\delta_{t}\bm{u}_{t}$; evaluate $f(\hat{\bm{x}}_{t})$; and collect the latest $m$ samples $\{(\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1}))\}_{i=1}^{m}$.
8:  Solve for the linear coefficient $\bm{g}_{t}$ via Eq. (8).
9:  Update $\bm{x}_{t+1}=\bm{x}_{t}-\eta\,\bm{g}_{t}$.
10:end for
@@
In particular, the following two key approaches are proposed to further improve the performance of Algorithm 1.
Adaptive Smoothing Radius. In Step 6, we propose an adaptive smoothing radius scheme (10) to enhance convergence instead of using a fixed $\delta$. The key idea is to adjust the smoothing radius $\delta_{t}$ in proportion to the magnitude of $\|\bm{g}_{t}\|$, since the update in Step 9 follows $\|\bm{x}_{t}\!-\!\bm{x}_{t-1}\|=\eta\|\bm{g}_{t}\|$. Using a fixed $\delta$ would cause convergence to stall at an error level on the order of $\mathcal{O}(\delta)$. In contrast, the adaptive scheme gradually reduces the perturbation size as the gradient norm decreases, enabling high convergence precision. See the proof of Lemma 1 in the Appendix for more details, where this adaptive scheme is used to control the error magnitude of the gradient estimator.